OpenStudy (anonymous):

P(t) = 1000+360sin((π/6)t)+180sin((π/3)t) Find the derivative and then solve for when P(t)=0

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[P'(t)=60\pi(cos(\frac{\pi}{6}t)+60\pi cos(\frac{\pi}{3}t)\]